How would I set up a system of equations for "Sue is older than Ken. The difference of their ages is 12 and the sum of their ages is 50." And how would I find the ages of each of them?
Answer by ankor@dixie-net.com(22740)

How would I set up a system of equations for "Sue is older than Ken.
The difference of their ages is 12 and the sum of their ages is 50."
And how would I find the ages of each of them?
:
Write an equation for each sentence
;
"The difference of their ages is 12"
s - k = 12
:
"the sum of their ages is 50."
s + k = 50
:
Use elimination
s - k = 12
s + k = 50
--------------adding eliminates k, find s
2s = 61
s = 31 is Sue's age, you can find Ken's age I'm sure